OpenStudy (anonymous):

What polynomial has roots of -6, 1, and 4?

hero (hero):

If a polynomial has roots of -6, 1, and 4, Then x = -6 x = 1 x = 4 Move everything to the left to get x + 6 = 0 x - 1 = 0 x - 4 = 0 Notice that x + 6, x - 1, x - 4 are factors of the polynomial By zero product property (x + 6)(x - 1)(x - 4) = 0 And f(x) = (x + 6)(x - 1)(x - 4) Expand the RHS to express the cubic poly in standard form.